# APCTP-Pre2023-010

### v2

@@ Abstract
We investigate the enhancement of the power spectra in models with multiple scalar fields large enough to produce primordial black holes. We present analytic solutions for the perturbations in the case of constant turns without the need for an effective field theory for the first time and clarify the role of the Hubble friction that has been overlooked previously. For an arbitrary number of fields we derive the criteria that can lead to an exponential amplification of the curvature perturbation on subhorizon scales, while leaving the perturbations stable on superhorizon scales. Finally, we apply our results on a three-field generalization of the “ultra-light” scenario and show how the presence of field-space torsion can yield distinct observables compared to the two-field case.
I Introduction
@@
The physical evolution of the observed universe is very well described by the hot big bang universe. Supported by solid observational evidences including the expansion of the universe, abundance of light elements, and the cosmic microwave background (CMB), the concordant cosmological model – called $\Lambda$CDM – specified by six parameters exhibits surprising agreement with the most recent observations on the CMB Planck:2018vyg . Yet, a successful hot big bang evolution demands a set of extremely finely tuned initial conditions. One representative example of such conditions is the homogeneity and isotropy of the CMB beyond the causal patches. These difficulties regarding the initial conditions can be resolved by introducing, before commencing the standard hot big bang, a period of accelerated expansion of the universe, called inflation Starobinsky:1980te ; Sato:1981ds ; Sato:1980yn ; Kazanas:1980tx ; Guth:1980zm ; Linde:1981mu ; Albrecht:1982wi . During inflation the expansion is so fast such that the whole observed universe was once in a single causally connected patch, which explains the homogeneity and isotropy of the CMB. Furthermore, quantum fluctuations during inflation are stretched and become the primordial perturbations, which seed the observable structure in the current universe. The properties of these primordial perturbations have been constrained by cosmological observation programmes with increasing accuracy during last decades, and are consistent with most recent data Planck:2018jri .
@@
While the general paradigm of inflation is robust, the realization of inflation is not an easily surmountable problem. To implement a period of inflation, we usually invoke a scalar field, called the inflaton, with a sufficiently flat potential energy that dominates the energy density of the universe due to the Hubble friction. Then the inflaton “slowly rolls” down the potential and behaves effectively as a cosmological constant, leading to the exponential expansion of the universe – inflation Mukhanov:2005sc ; Weinberg:2008zzc ; Lyth:2009zz . Under this slow-roll approximation, the calculations of the two-point correlation functions of the primordial perturbations, or the power spectra, are straightforward and we can constrain easily the functional form of the inflaton potential by data from the anisotropies of the CMB and the large-scale inhomogeneous distribution of galaxies Lidsey:1995np .
@@
However, in the standard model of particle physics (SM) that has been verified and confirmed experimentally, there is no fundamental scalar field that can play the role of the inflaton. [^1] Note that by introducing a non-minimal coupling to gravity, the SM Higgs field may play the role of the inflaton Bezrukov:2007ep . Thus, the realization of inflation resort to the theories beyond SM, or in the context of effective field theory. One common feature of these theories is the existence of degrees of freedom with heavier mass scales than our interest. Since it is very likely that the energy scale of the early universe is far higher than that proved by terrestrial particle accelerators, as high as $10^{15}$ GeV, those heavy degrees of freedom which play little role in the SM could have actively participated in the inflationary dynamics. Then, these participating degrees of freedom are relevant for inflation and thus can all be identified as the inflatons. Therefore there are good motivations to consider multiple number of degrees of freedom during inflation – multi-field inflation. For reviews see e.g. Bassett:2005xm ; Wands:2007bd ; Langlois:2010xc ; Gong:2016qmq .
@@
Furthermore, recent cosmological observations have been bringing tantalizing phenomenological hints indicating that we need more than the six-parameter standard cosmological model. There are a number of outliers in the power spectrum of the temperature fluctuations of the CMB, which could be accommodated by “features” in the inflationary dynamics caused by various reasons Chluba:2015bqa . The rotation angle of the CMB polarizations is found to be non-zero Minami:2020odp ; Eskilt:2022cff , which calls for possible parity-violating mechanism coupled to the photons such as the Chern-Simons term with axion Carroll:1989vb . The value of the Hubble parameter determined by the local measurements including supernovae is not in agreement with the value determined by the CMB with very high significance Riess:2019cxk . Currently there is no leading candidate to resolve this discrepancy, but many possibilities have been suggested including e.g. early domination of dark energy Poulin:2018cxd . All these observations indicate that the standard cosmological model needs to be refined, including the constituents yet to be discovered.
@@
Especially, the discovery of the gravitational waves (GWs) by LIGO LIGOScientific:2016aoc has spurred many studies on various aspects of GWs and related topics. A very interesting possibility is that the binary black holes relevant for the merger events are not the remnants of the stellar evolution but of primordial origin – primordial black holes (PBHs) Zeldovich:1967lct ; Hawking:1971ei ; Carr:1974nx . It is believed that PBHs are formed when the amplitude of the primordial perturbations at certain length scale is so high that as soon as such a region with very high perturbation amplitude enters the horizon after inflation, gravitational collapse immediately follows and the corresponding horizon-size black hole is formed. Therefore, for the PBHs, we need a large enhancement of the power spectrum compared to the value constrained on the CMB scales, ${\cal O}(10^{-5})$. The exact number necessary for the generation of PBHs depends on the details of the numerical studies, but generally it is estimated that the value should be ${\cal O}(10^{-1}-10^{-2})$ Harada:2013epa . Thus, we need an amplification of the primordial perturbations by ${\cal O}(10^{3}-10^{4})$ around the scale that corresponds to the size, or equivalently the mass, of the PBHs.
@@
Such a large enhancement in single-field inflation is, however, not arbitrary but is limited with the steepest growth of the power spectrum being $k^{4}$ Byrnes:2018txb ; Carrilho:2019oqg . The resulting PBH mass and the prospects for LIGO and future observations are highly constrained. In multi-field inflation, the amplification of the primordial perturbations has also been discussed in many different contexts, e.g. oscillations in the inflationary trajectory Pi:2017gih , hybrid-inspired models Garcia-Bellido:1996mdl ; Cheong:2019vzl ; Braglia:2020eai ; Braglia:2022phb ; Cheong:2022gfc , (sudden) turns in field space Palma:2020ejf ; Fumagalli:2020nvq ; Fumagalli:2020adf ; Kolb:2022eyn ; Anguelova:2020nzl ; Bhattacharya:2022fze ; Fumagalli:2021mpc ; Aragam:2023adu , resonance in the speed of sound Cai:2018tuh , multiple features Boutivas:2022qtl to list only a few examples.
@@
In this article, we investigate the general conditions for the enhancement of the power spectra of perturbations during multi-field inflation. This article is outlined as follows. In section II, we first provide a concise review of multi-field inflation to introduce our notations and necessary materials for the following discussions. In section III we derive analytic solution for the enhancement of the curvature perturbation in two-field models that undergo constant turns. In section IV we describe the conditions under which perturbations for an arbitrary number of fields may experience a temporary exponential enhancement in the subhorizon regime but remain under control on superhorizon scales. In section V we investigate the models with isometries and focus on ultra-light isocurvature perturbations. Finally, in section VI we summarize our study.
II Multi-field perturbations
@@
The purpose of this section, devoted for a short and self-contained review, is four-fold. First, we introduce the notations we will adopt throughout this article. Second, we set up the Lagrangian to be analyzed in the next sections. Third, we discuss some idea relevant for following contents, and lastly we adopt the Hamiltonian formulation of inflationary perturbations.
II.1 Orthonormal frame and initial conditions
@@
$S=\int d^{4}x\sqrt{-g}\bigg{[}\frac{R}{2}-\frac{1}{2}G_{IJ}g^{\mu\nu}\partial_{\mu}\phi^{I}\partial_{\nu}\phi^{J}-V(\phi)\bigg{]}\,.$ (1)
@@
$Q^{I}\equiv\delta\phi^{I}+{\psi\over H}\dot{\phi}^{I}\,,$ (2)
is given by the following expression Gong:2011uw :
$S_{2}={1\over 2}\int\,\mathrm{d}t\,\mathrm{d}^{3}xa^{3}\left(\,\text{D}_{t}Q^{I}\,\text{D}_{t}Q_{I}-{1\over a^{2}}\nabla Q_{I}\nabla Q^{I}-M_{IJ}Q^{I}Q^{J}\right)\,,$ (3)
where the field index is raised and lowered by the field space metric $G_{IJ}$, and the mass matrix $M_{IJ}$ is defined as follows Achucarro:2010da :
$M_{IJ}$ $\equiv V_{;IJ}-R_{IKLJ}\dot{\phi}^{K}\dot{\phi}^{L}+(3-\epsilon)\dot{\phi}_{I}\dot{\phi}_{J}+{1\over H}\left(\dot{\phi}_{I}V_{,J}+\dot{\phi}_{J}V_{,I}\right)\,,$ (4)
@@
$\,\text{D}_{t}A^{I}\equiv{\,\mathrm{d}\over\,\mathrm{d}t}A^{I}+\Gamma^{I}_{JK}\dot{\phi}^{K}A^{J}\,,$ (5)
@@
$R^{I}_{KLJ}\equiv\partial_{L}\Gamma^{I}_{KJ}-\partial_{J}\Gamma^{I}_{KL}+\Gamma^{I}_{LN}\Gamma^{N}_{KJ}-\Gamma^{I}_{JN}\Gamma^{N}_{KL}\,.$ (6)
Recall that the field fluctuations behave as vectors in the field space Gong:2011uw and the use of the covariant derivative operator $\,\text{D}_{t}$ ensures that the action is covariant under arbitrary field redefinitions.
For computational ease we change to a local orthonormal frame  Achucarro:2010da . This can be done by defining a new set of orthonormal basis vectors $\{\hat{\bm{e}}_{a}\}$ according to
$\begin{split}\hat{\bm{e}}_{a}&\equiv E^{~{}I}_{a}\bm{e}_{I}\,,\\ \bm{e}_{I}&\equiv\tilde{E}^{~{}a}_{I}\hat{\bm{e}}_{a}\,,\end{split}$ (7)
@@
$\begin{split}E^{~{}I}_{a}E_{b}^{~{}J}G_{IJ}&=\delta_{ab}\,,\\ \tilde{E}^{~{}a}_{I}\tilde{E}^{~{}b}_{J}\delta_{ab}&=G_{IJ}\,,\\ G_{IJ}E^{~{}J}_{a}&=\delta_{ab}\tilde{E}^{~{}b}_{I}\,.\end{split}$ (8)
The last equation relates the components of $\tilde{E}$ with the metric tensor and $E$. This new set of basis vectors constitutes a non-coordinate basis, as the commutator between those basis vectors does not vanish: $[\hat{\bm{e}}_{a},\hat{\bm{e}}_{b}]\neq 0$. The next step is to define a covariant derivative in the new frame. Taking the covariant time derivative of (7) gives
$\,\text{D}_{t}\hat{\bm{e}}_{a}=\left(\dot{E}^{~{}I}_{a}+E^{~{}J}_{a}\Gamma^{I}_{JK}\dot{\phi}^{K}\right)\tilde{E}^{~{}b}_{I}\hat{\bm{e}}_{b}\,,$ (9)
@@
$z_{a}{}^{b}\equiv\left(\dot{E}^{~{}I}_{a}+E^{~{}J}_{a}\Gamma^{I}_{JK}\dot{\phi}^{K}\right)\tilde{E}^{~{}b}_{I}\,,$ (10)
which is antisymmetric in its two indices [^2] Note that for tensors defined in the orthonormal base, we raise and lower indices using the Euclidean metric $\delta_{ab}$ so there is no distinction between the two indices. However, we keep track on the position of indices in each column. and can be used to define a connection with respect to the non-coordinate basis:
$\,\text{D}_{t}\hat{\bm{e}}_{a}=z^{~{}b}_{a}\hat{\bm{e}}_{b}\,.$ (11)
Until now, our discussions on the change to a set of orthonormal basis is completely general. From now on, we identify the orthonormal basis as the kinematic one so that the first two basis vectors are tangent and normal to the background trajectory Gordon:2000hv ; Achucarro:2012sm . The antisymmetric matrix $z$ has non-zero components only above and lower of the principal diagonal, which allows us to write the previous system of equations as
$\,\text{D}_{t}\begin{pmatrix}\hat{\bm{e}}_{\rm t}\\ \hat{\bm{e}}_{\rm n}\\ \hat{\bm{e}}_{\rm b}\\ \vdots\end{pmatrix}=\begin{pmatrix}0&\omega&0&\cdots\\ -\omega&0&\tau_{\rm b}&\cdots\\ 0&-\tau_{\rm b}&0&\cdots\\ \vdots&\vdots&\vdots&\ddots\end{pmatrix}\begin{pmatrix}\hat{\bm{e}}_{\rm t}\\ \hat{\bm{e}}_{\rm n}\\ \hat{\bm{e}}_{\rm b}\\ \vdots\end{pmatrix}\,,$ (12)
also known as the Frenet-Serret system kreyszig2013differential . The first three basis vectors are known as the tangent, normal and binormal vectors and the parameters $\omega$ and $\tau_{\rm b}$ are the curvature (commonly known as the turn rate in the inflationary literature) and the torsion respectively. In the orthonormal basis the components of perturbations are expressed as
$Q^{I}=E^{~{}I}_{a}q^{a}\,,$ (13)
@@
$\,\text{D}_{t}\boldsymbol{q}=\,\text{D}_{t}(q^{a}\hat{\bm{e}}_{a})=\left(\dot{q}^{a}+z^{~{}a}_{b}q^{b}\right)\hat{\bm{e}}_{a}=\left(\dot{q}^{a}-z^{a}_{~{}b}q^{b}\right)\hat{\bm{e}}_{a}\,,$ (14)
@@
$S_{2}={1\over 2}\int\,\mathrm{d}t\,\mathrm{d}^{3}xa^{3}\left[\delta_{ab}\dot{q}^{a}\dot{q}^{b}-2z_{ab}\dot{q}^{a}q^{b}-{1\over a^{2}}\delta_{ab}\nabla q^{a}\nabla q^{b}-\left(M_{IJ}E^{~{}I}_{a}E^{~{}J}_{b}+z_{ac}z^{c}_{~{}b}\right)q^{a}q^{b}\right]\,,$ (15)

### v3

@@ Abstract
We investigate the enhancement of the power spectra large enough to produce primordial black holes in models with multiple scalar fields. We present analytic solutions for the perturbations in the case of constant turns without the need for an effective field theory for the first time and clarify the role of the Hubble friction that has been overlooked previously. We derive the criteria for an arbitrary number of fields that can lead to an exponential amplification of the curvature perturbation on subhorizon scales, while leaving the perturbations stable on superhorizon scales. Finally, we apply our results to a three-field generalization of the “ultra-light” scenario and show how the presence of field-space torsion can yield distinct observables compared to the two-field case.
1 Introduction
@@
The physical evolution of the observed universe is very well described by the hot big bang universe. Supported by solid observational evidences including the expansion of the universe, abundance of light elements, and the cosmic microwave background (CMB), the concordant cosmological model – called $\Lambda$CDM – specified by six parameters exhibits surprising agreement with the most recent observations on the CMB [1]. Yet, a successful hot big bang evolution demands a set of extremely finely tuned initial conditions. One representative example of such conditions is the homogeneity and isotropy of the CMB beyond the causal patches. These difficulties regarding the initial conditions can be resolved by introducing, before commencing the standard hot big bang, a period of accelerated expansion of the universe, called inflation [2, 3, 4, 5, 6, 7, 8]. During inflation the expansion is so fast such that the whole observed universe was once in a single causally connected patch, which explains the homogeneity and isotropy of the CMB. Furthermore, quantum fluctuations during inflation are stretched and become the primordial perturbations, which seed the observable structure in the current universe. The properties of these primordial perturbations have been constrained by cosmological observation programmes with increasing accuracy during the last decades, and are consistent with most recent data [9].
@@
While the general paradigm of inflation is robust, the realization of inflation is not an easily surmountable problem. To implement a period of inflation, we usually invoke a scalar field, called the inflaton, with a sufficiently flat potential energy that dominates the energy density of the universe due to the Hubble friction. Then the inflaton “slowly rolls” down the potential and behaves effectively as a cosmological constant, leading to the exponential expansion of the universe – inflation [10, 11, 12]. Under this slow-roll approximation, the calculations of the two-point correlation functions of the primordial perturbations, or the power spectra, are straightforward and we can easily constrain the functional form of the inflaton potential by data from the anisotropies of the CMB and the large-scale inhomogeneous distribution of galaxies [13].
@@
However, in the standard model of particle physics (SM) that has been verified and confirmed experimentally, there is no fundamental scalar field that can play the role of the inflaton. [^1] Note that by introducing a non-minimal coupling to gravity, the SM Higgs field may play the role of the inflaton [14]. Thus, the realization of inflation resorts to the theories beyond SM, or in the context of effective field theory. One common feature of these theories is the existence of degrees of freedom with heavier mass scales than our interest. Since it is very likely that the energy scale of the early universe is far higher than that proved by terrestrial particle accelerators, as high as $10^{15}$ GeV, those heavy degrees of freedom which play little role in the SM could have actively participated in the inflationary dynamics. Then, these participating degrees of freedom are relevant for inflation and thus can all be identified as the inflatons. Therefore there are good motivations to consider multiple degrees of freedom during inflation – multi-field inflation. For reviews see e.g. [15, 16, 17, 18].
@@
Furthermore, recent cosmological observations have brought tantalizing phenomenological hints indicating that we need more than the six-parameter standard cosmological model. There are several outliers in the power spectrum of the temperature fluctuations of the CMB, which could be accommodated by “features” in the inflationary dynamics caused by various reasons [19]. The rotation angle of the CMB polarizations is found to be non-zero [20, 21], which calls for possible parity-violating mechanism coupled to the photons such as the Chern-Simons term with axion [22]. The value of the Hubble parameter determined by the local measurements including supernovae is not in agreement with the value determined by the CMB with very high significance [23]. Currently there is no leading candidate to resolve this discrepancy, but many possibilities have been suggested including e.g. early domination of dark energy [24]. All these observations indicate that the standard cosmological model needs to be refined, including the constituents yet to be discovered.
@@
Especially, the discovery of gravitational waves (GWs) by LIGO [25] has spurred many studies on various aspects of GWs and related topics. An exciting possibility is that the binary black holes relevant for the merger events are not the remnants of the stellar evolution but of primordial origin – primordial black holes (PBHs) [26, 27, 28]. It is believed that PBHs are formed when the amplitude of the primordial perturbations at certain length scale is so high that as soon as such a region with very high perturbation amplitude enters the horizon after inflation, gravitational collapse immediately follows and the corresponding horizon-size black hole is formed. Therefore, for the PBHs, we need a large enhancement of the power spectrum compared to the value constrained on the CMB scales, ${\cal O}(10^{-5})$. The exact number necessary for the generation of PBHs depends on the details of the numerical studies, but generally it is estimated that the value should be ${\cal O}(10^{-1}-10^{-2})$ [29]. Thus, we need an amplification of the primordial perturbations by ${\cal O}(10^{3}-10^{4})$ around the scale that corresponds to the size, or equivalently the mass, of the PBHs.
@@
Such a large enhancement in single-field inflation is, however, not arbitrary but is limited with the steepest growth of the power spectrum being $k^{4}$ [30, 31]. The resulting PBH mass and the prospects for LIGO and future observations are highly constrained. In multi-field inflation, the amplification of the primordial perturbations has also been discussed in many different contexts, e.g. oscillations in the inflationary trajectory [32], hybrid-inspired models [33, 34, 35, 36, 37], (sudden) turns in field space [38, 39, 40, 41, 42, 43, 44, 45], resonance in the speed of sound [46], multiple features [47] to list only a few examples.
@@
In this article, we investigate the general conditions for the enhancement of the power spectra of perturbations during multi-field inflation. This article is outlined as follows. In section 2, we first provide a concise review of multi-field inflation to introduce our notations and necessary materials for the following discussions. In section 3 we derive analytic solution for the enhancement of the curvature perturbation in two-field models that undergo constant turns. In section 4 we describe the conditions under which perturbations for an arbitrary number of fields may experience a temporary exponential enhancement in the subhorizon regime but remain under control on superhorizon scales. In section 5 we investigate the models with isometries and focus on ultra-light isocurvature perturbations. Finally, in section 6 we summarize our study.
2 Multi-field perturbations
@@
The purpose of this section, devoted to a short and self-contained review, is four-fold. First, we introduce the notations we will adopt throughout this article. Second, we set up the Lagrangian to be analyzed in the next sections. Third, we discuss some ideas relevant to the following contents, and lastly we adopt the Hamiltonian formulation of inflationary perturbations.
2.1 Orthonormal frame and initial conditions
@@
$S=\int d^{4}x\sqrt{-g}\bigg{[}\frac{R}{2}-\frac{1}{2}G_{IJ}g^{\mu\nu}\partial_{\mu}\phi^{I}\partial_{\nu}\phi^{J}-V(\phi)\bigg{]}\,.$ (2.1)
@@
$Q^{I}\equiv\delta\phi^{I}+{\psi\over H}\dot{\phi}^{I}\,,$ (2.2)
is given by the following expression [48]:
$S_{2}={1\over 2}\int\,\mathrm{d}t\,\mathrm{d}^{3}xa^{3}\left(\,\text{D}_{t}Q^{I}\,\text{D}_{t}Q_{I}-{1\over a^{2}}\nabla Q_{I}\nabla Q^{I}-M_{IJ}Q^{I}Q^{J}\right)\,,$ (2.3)
where the field index is raised and lowered by the field space metric $G_{IJ}$, and the mass matrix $M_{IJ}$ is defined as follows [49]:
$M_{IJ}$ $\equiv V_{;IJ}-R_{IKLJ}\dot{\phi}^{K}\dot{\phi}^{L}+(3-\epsilon)\dot{\phi}_{I}\dot{\phi}_{J}+{1\over H}\left(\dot{\phi}_{I}V_{,J}+\dot{\phi}_{J}V_{,I}\right)\,,$ (2.4)
@@
$\,\text{D}_{t}A^{I}\equiv{\,\mathrm{d}\over\,\mathrm{d}t}A^{I}+\Gamma^{I}_{JK}\dot{\phi}^{K}A^{J}\,,$ (2.5)
@@
$R^{I}_{KLJ}\equiv\partial_{L}\Gamma^{I}_{KJ}-\partial_{J}\Gamma^{I}_{KL}+\Gamma^{I}_{LN}\Gamma^{N}_{KJ}-\Gamma^{I}_{JN}\Gamma^{N}_{KL}\,.$ (2.6)
Recall that the field fluctuations behave as vectors in the field space [48] and the use of the covariant derivative operator $\,\text{D}_{t}$ ensures that the action is covariant under arbitrary field redefinitions.
For computational ease we change to a local orthonormal frame  [49]. This can be done by defining a new set of orthonormal basis vectors $\{\hat{\bm{e}}_{a}\}$ according to
$\begin{split}\hat{\bm{e}}_{a}&\equiv E^{~{}I}_{a}\bm{e}_{I}\,,\\ \bm{e}_{I}&\equiv\tilde{E}^{~{}a}_{I}\hat{\bm{e}}_{a}\,,\end{split}$ (2.7)
@@
$\begin{split}E^{~{}I}_{a}E_{b}^{~{}J}G_{IJ}&=\delta_{ab}\,,\\ \tilde{E}^{~{}a}_{I}\tilde{E}^{~{}b}_{J}\delta_{ab}&=G_{IJ}\,,\\ G_{IJ}E^{~{}J}_{a}&=\delta_{ab}\tilde{E}^{~{}b}_{I}\,.\end{split}$ (2.8)
The last equation relates the components of $\tilde{E}$ with the metric tensor and $E$. This new set of basis vectors constitutes a non-coordinate basis, as the commutator between those basis vectors does not vanish: $[\hat{\bm{e}}_{a},\hat{\bm{e}}_{b}]\neq 0$. The next step is to define a covariant derivative in the new frame. Taking the covariant time derivative of (2.7) gives
$\,\text{D}_{t}\hat{\bm{e}}_{a}=\left(\dot{E}^{~{}I}_{a}+E^{~{}J}_{a}\Gamma^{I}_{JK}\dot{\phi}^{K}\right)\tilde{E}^{~{}b}_{I}\hat{\bm{e}}_{b}\,,$ (2.9)
@@
$z_{a}{}^{b}\equiv\left(\dot{E}^{~{}I}_{a}+E^{~{}J}_{a}\Gamma^{I}_{JK}\dot{\phi}^{K}\right)\tilde{E}^{~{}b}_{I}\,,$ (2.10)
which is antisymmetric in its two indices [^2] Note that for tensors defined in the orthonormal base, we raise and lower indices using the Euclidean metric $\delta_{ab}$ so there is no distinction between the two indices. However, we keep track of the position of indices in each column. and can be used to define a connection with respect to the non-coordinate basis:
$\,\text{D}_{t}\hat{\bm{e}}_{a}=z^{~{}b}_{a}\hat{\bm{e}}_{b}\,.$ (2.11)
Until now, our discussions on the change to a set of orthonormal basis is completely general. From now on, we identify the orthonormal basis as the kinematic one so that the first two basis vectors are tangent and normal to the background trajectory [50, 51]. The antisymmetric matrix $z$ has non-zero components only above and lower of the principal diagonal, which allows us to write the previous system of equations as
$\,\text{D}_{t}\begin{pmatrix}\hat{\bm{e}}_{\rm t}\\ \hat{\bm{e}}_{\rm n}\\ \hat{\bm{e}}_{\rm b}\\ \vdots\end{pmatrix}=\begin{pmatrix}0&\omega&0&\cdots\\ -\omega&0&\tau_{\rm b}&\cdots\\ 0&-\tau_{\rm b}&0&\cdots\\ \vdots&\vdots&\vdots&\ddots\end{pmatrix}\begin{pmatrix}\hat{\bm{e}}_{\rm t}\\ \hat{\bm{e}}_{\rm n}\\ \hat{\bm{e}}_{\rm b}\\ \vdots\end{pmatrix}\,,$ (2.12)
also known as the Frenet-Serret system [52]. The first three basis vectors are known as the tangent, normal and binormal vectors and the parameters $\omega$ and $\tau_{\rm b}$ are the curvature (commonly known as the turn rate in the inflationary literature) and the torsion respectively. In the orthonormal basis the components of perturbations are expressed as
$Q^{I}=E^{~{}I}_{a}q^{a}\,,$ (2.13)
@@
$\,\text{D}_{t}\boldsymbol{q}=\,\text{D}_{t}(q^{a}\hat{\bm{e}}_{a})=\left(\dot{q}^{a}+z^{~{}a}_{b}q^{b}\right)\hat{\bm{e}}_{a}=\left(\dot{q}^{a}-z^{a}_{~{}b}q^{b}\right)\hat{\bm{e}}_{a}\,,$ (2.14)
@@
$S_{2}={1\over 2}\int\,\mathrm{d}t\,\mathrm{d}^{3}xa^{3}\left[\delta_{ab}\dot{q}^{a}\dot{q}^{b}-2z_{ab}\dot{q}^{a}q^{b}-{1\over a^{2}}\delta_{ab}\nabla q^{a}\nabla q^{b}-\left(M_{IJ}E^{~{}I}_{a}E^{~{}J}_{b}+z_{ac}z^{c}_{~{}b}\right)q^{a}q^{b}\right]\,,$ (2.15)